OpenStudy (anonymous):

"Find the exact value of cos(θ + β) given that sinθ = 7/25 and cosβ = 12/13. (Both θ and β are in Quadrant II)" I got as far as setting it up as 'Cos (θ + β) = cos(-24/25) cos(-12/13) - sin(7/25) sin(5/13)' but now I'm stuck and don't know how to go forward with it.

OpenStudy (anonymous):

oops, a common mistake

OpenStudy (anonymous):

it is just \[\frac{-24}{25}\times \frac{-12}{13}-\frac{7}{25}\times \frac{5}{13}\]

OpenStudy (anonymous):

Ok, so I just multiply the two cos fractions, then the sin fractions, and finally find the difference?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

don't take the cosine OF -12/13, the cosine IS -12/13 etc